Consider the following tasks that must be assigned to four workers on a conveyor--paced assembly line (i.e., a machine--‐paced line flow). Each worker must perform at least one task. Time to Complete Task (seconds): Task 1: 26 Task 2: 25 Task 3: 15 Task 4: 24 Task 5: 15 Task 6: 20 Task 7: 47 Task 8: 15 Task 9: 20 Task 10:25 Task 11: 15 Task 12: 20 The current conveyor--paced assembly-‐line configuration assigns the workers in the following way: • Worker 1: Tasks1, 2, 3 • Worker 2: Tasks 4, 5, 6 • Worker 3: Tasks 7, 8, 9 • Worker 4: Tasks 10, 11, 12 Assume a demand rate of 21 units per minute. What would be the target manpower?
Answer #1

Total task time = sum of all task times =

Total Task Time = 26+25+15+24+15+20+47+15+20+25+15+20

Total Task time = 267 seconds

Desired output = 21 units per minutes

Cycle time or takt time = Available time / desired output = 60/21 = 2.857143 seconds per unit.

Target manpower = Total task time / cycle time =

Target Manpower = 267 / 2.857143 = 93.45

Target Manpower = 94